# Optimal Investment Portfolio of Thyristor- and IGBT-based Electrolysis Rectifiers in Utility-scale Renewable P2H Systems

### v1

@@ Abstract
Renewable power-to-hydrogen (ReP2H) systems require rectifiers to supply power to electrolyzers (ELZs). Two main types of rectifiers, insulated-gate bipolar transistor rectifiers (IGBT-Rs) and thyristor rectifiers (TRs), offer distinct tradeoffs. IGBT-Rs provide flexible reactive power control but are costly, whereas TRs are more affordable with lower power loss but consume a large amount of uncontrollable reactive power. A mixed configuration of rectifiers in utility-scale ReP2H systems could achieve an decent tradeoff and increase overall profitability. To explore this potential, this paper proposes an optimal investment portfolio model. First, we model and compare the active and reactive power characteristics of ELZs powered by TRs and IGBT-Rs. Second, we consider the investment of ELZs, rectifiers, and var resources and coordinate the operation of renewables, energy storage, var resources, and the on-off switching and load allocation of multiple ELZs. Subsequently, a two-stage stochastic programming (SP) model based on weighted information gap decision theory (W-IGDT) is developed to address the uncertainties of the renewable power and hydrogen price, and we apply the progressive hedging (PH) algorithm to accelerate its solution. Case studies demonstrate that optimal rectifier configurations increase revenue by at most 2.56% compared with using only TRs or IGBT-Rs, as well as those in existing projects. Under the optimal portfolio, reactive power compensation investment is nearly eliminated, with a preferred TR-to-IGBT-R ratio of 3:1.
Index Terms:
renewable power-to-hydrogen, electrolyzer, rectifier, investment portfolio, reactive power, stochastic programmingNomenclature
@@
Active/reactive power of the ELZ
Active power of the rectifier of the ELZ
@@ I-A Background and Motivation
Renewable power-to-hydrogen (ReP2H) facilitates wind and solar integration [1] and has significant potential for applications in energy storage, transportation [1], and the chemical industry [2]. With the continuous growth of renewable energy installations, the ReP2H industry has grown rapidly under policy incentives and market drivers [3, 4], leading to increased investment in electrolyzers (ELZs) in large-scale hydrogen plants to replace fossil-based hydrogen production.
Industrial hydrogen plants typically consist of multiple ELZs [5], each integrating a rectifier, an electrolytic stack, and the balance of plant (BoP) [6]. Since electrolytic stacks operate on DC power, rectifiers are required to convert AC to DC. Currently, two types of rectifiers are commercially available: thyristor rectifiers (TRs) and insulated-gate bipolar transistor-based rectifiers (IGBT-Rs) [7]. They offer distinct tradeoffs, as summarized in Table I.
IGBT-Rs use PWM modulation and voltage vector control, enabling independent active and reactive power control. Thus, IGBT-Rs can maintain a power factor (PF) of 1 on the AC side [8] or provide flexible reactive power support [9], or even enable voltage ride-through [10]. Moreover, their fast response makes IGBT-Rs well suited for weak or off-grid systems [11]. However, their structural complexity leads to higher costs and power losses.
@@
In practice, some large-scale projects [13, 14] use mixed rectifier (MR) configurations to leverage the complementary characteristics of TRs and IGBT-Rs. The China Energy Engineering Songyuan Hydrogen Industry Park [13] in Songyuan, China, employs 12 TRs and 24 IGBT-Rs to power 36 alkaline ELZs, and the Wind-Solar Hydrogen-Ammonia Integration Project [14] in Da’an, China, uses 32 TRs and 32 IGBT-Rs for 64 ELZs. However, to the best of the author’s knowledge, such configurations have not been quantitatively analyzed. Suboptimal configurations may lead to excessive or insufficient reactive power compensation, increasing investment or operating costs.
Moreover, our previous research [6, 15] revealed that TR-powered ELZs face tradeoffs between efficiency and the reactive load. This means that the investment and operation of the ReP2H system should be coordinated in terms of both active and reactive power. However, existing ReP2H planning research has focused mainly on the sizing of energy sources and ELZs while considering only active power and neglecting rectifier selection and var resources. An optimal investment decision from the perspective of active and reactive power coordination is vital for improving capital efficiency, which, unfortunately, has not been investigated.
@@ I-B Literature Review and Research Gap
ReP2H planning research typically addresses renewable energy, battery energy storage (BES), ELZs, and hydrogen storage (HS). For example, Li et al. [18] proposed an investment model for ELZs and seasonal HS, considering power and hydrogen delivery, to address the spatiotemporal imbalance between renewable resources and hydrogen demand. Pan et al. [19] developed a robust capacity sizing model for wind turbines (WTs), photovoltaics (PVs), ELZs, and HSs in integrated energy systems. Zhu et al. [20] proposed an energy management framework and studied BES capacity sizing in off-grid ReP2H systems. Li et al. [21] explored the configuration of the number and size of ELZs and developed operation rules to enhance wind power integration. To assess the cost competitiveness of green hydrogen, Ibáñez-Rioja et al. [22] optimized planning and operation using 30 years of real-world wind and solar data. Similarly, Zheng et al. [23] analyzed the levelized cost of hydrogen (LCOH) for off-grid wind P2H systems and reported that optimal sizing reduces LCOH.
However, the aforementioned studies [18, 19, 20, 21, 22, 23] overlooked rectifier configuration and reactive power balance in ReP2H planning. Given that rectifiers account for 20–30% of ELZ investment [24], their optimal selection is crucial for economic feasibility. Current research on rectifiers for hydrogen production has focused mainly on performance analysis [12, 8] or short-term scheduling [6, 15].
For example, Ruuskanen et al. [12] investigated the power quality and reactive power characteristics of ELZs powered by TRs. Koponen et al. [8] compared the tradeoffs of the TR and IGBT-R in terms of performance and investment in industrial applications. In our prior work [15, 6], Li et al. [15] integrated rectifier power characteristics into ReP2H scheduling, revealing conflicts between productivity and the grid-side PF in grid-connected hydrogen plants, whereas Zeng et al. [6] addressed the tradeoffs between P2H energy conversion efficiency and network losses caused by the reactive power of a TR and proposed a coordinated power management approach for large hydrogen plants that integrate multiple ELZs.
@@ I-C Contributions of This Work
The active and reactive characteristics of TR- and IGBT-R-powered ELZs are modeled and compared, revealing their complementarity and the potential of a mixed rectifier configuration, which is incorporated into planning.
@@
An optimal investment portfolio model for MR and var resources in ReP2H systems is proposed. The model handles uncertainties via weighted information gap decision theory (W-IGDT). A two-stage decomposition algorithm, progressive hedging (PH) [25], is applied to improve computational efficiency.
@@
Case studies on realistic systems show that optimizing MR configuration can increase revenue by at most 2.56% compared with that of uniform configurations or those in existing projects. The results suggest that a 3:1 ratio of TRs to IGBT-Rs offers better benefits.
@@ II-A General Configuration of ReP2H Systems
On the basis of existing projects [13, 14], the general structure of an ReP2H system is illustrated in Fig. 1. The system integrates renewable energy sources (wind and solar), transmission lines, a hydrogen plant, and sometimes a BES. The hydrogen plant often comprises multiple ELZs. Owing to policy restrictions [26], power exchange with the external grid is limited and thus is not considered in this work.
@@ II-C Reactive Power Characteristics
where is the turn ratio of the rectifier transformer and where is the harmonic factor of the 24-pulse TR. Since and depend on and , substituting (3) and (5) into (18) yields the complete expression for , which is used for MR configuration planning in Section III.
@@ II-C2 IGBT-R
II-C3 Comparison
@@
To illustrate the impact of different rectifier configurations, a straightforward example is given. A hydrogen plant with four 5 MW ELZs supplied with 12 MW of power for one hour is considered. Using hydrogen production maximization as the objective, we calculate the hydrogen yield and reactive power demand for each configuration. Fig. 5 presents the results along with the investment costs. TRs offer lower costs and slightly higher hydrogen production, whereas IGBT-Rs provide reactive power support, which is critical for voltage security, reducing AC network losses, lowering var compensator costs, and enhancing the overall profitability of the ReP2H system, as demonstrated in our previous work [6]. An MR configuration may achieve an optimal balance among these factors.
Remark 1.
Although the active and reactive powers of the TR are coupled, they are both governed by the electrolytic load and temperature. Managing the load allocation and temperature of multiple TR-powered ELZs can optimize the reactive power flow in the ReP2H system to improve overall productivity; see the tradeoff emphasized in Section II-E of [6]. In other words, the investment planning and production scheduling of the hydrogen plant should be coordinated with the electrical network.
@@ III Optimal Investment Portfolio Planning of Mixed Rectifier Configurations
III-A Planning Framework
| Operational constraints | |
| Power flow [30] | (21) |
| WTs and PV plants [6] | (29) |
| BES [31] | (31) |
| Var compensation | (35) |
| OLTC [6] | (36) |
The planning framework consists of two stages:
@@ III-B Objective Function
Here, and are the discount rate and equipment lifespan, respectively; represents the probability of scenario ; and is the annual O&M cost factor.
@@ III-C2 Operational constraints
We consider the coordination among active and reactive resources such as the electrical network, ELZs, WTs, PV plants, BES, var compensation, and OLTC. Their models are presented in Table III, where use the DistFlow model [30] to describe the network power flow, which is radial in common ReP2H systems.
In operation, active power balance is achieved by on-off switching and load allocation among multiple ELZs in the hydrogen plant, assisted by the BES. The reactive power flow is optimized in coordination with the active power to support the voltage and reduce network losses. The OLTC regulates the voltage at the hydrogen plant to ensure that the TR firing angle stays within a preferred range. For details of the operation model, readers are referred to [6].
@@ III-D2 W-IGDT-based portfolio model
To handle the intraday uncertainty of renewable power, a risk-averse model based on robust IGDT [34] is developed. Based on the revenue of (46), with a preset deviation factor , the robust IGDT model maximizes the uncertainty boundary under the worst-case renewable generation, as follows:
@@ IV Case Studies
Case studies are performed on two real-world systems in Inner Mongolia, China, as shown in Fig. 7. The test systems are adapted from our prior work [6]. The small-scale system validates the proposed portfolio model and explores the potential of MR configurations, whereas the large-scale system verifies the scalability of the model and provides insights for practice. Simulations are performed via Wolfram Mathematica 14.0, with optimizations solved via Gurobi 11.0.0.
@@ IV-A2 Large-scale system
| ELZ w/o rectifier | 5 MW | 7,000,000 CNY | 8% | 20 years |
| TR | 6 MVA | 1,000,000 CNY | ||
| IGBT-R | 6 MVA | 2,200,000 CNY | ||
| SVG | 350 CNY/kVar |
The key investment parameters are listed in Table IV. The topology and operational details of the ReP2H system and ELZs are available in [6]. Renewable generation for typical days, derived from on-site historical data via K-means clustering, is shown in Figs. 8(a) and 8(b). The hydrogen price scenario is shown in Fig. 8(c), with fluctuations modeled based on local agricultural ammonia demand [32].

### v2

@@ Abstract
Renewable power-to-hydrogen (ReP2H) systems require rectifiers to supply power to electrolyzers (ELZs). Two main types of rectifiers, insulated-gate bipolar transistor rectifiers (IGBT-Rs) and thyristor rectifiers (TRs), offer distinct tradeoffs. IGBT-Rs provide flexible reactive power control but are costly, whereas TRs are more affordable with lower power loss but consume a large amount of uncontrollable reactive power. A mixed configuration of rectifiers in utility-scale ReP2H systems could achieve a decent tradeoff and increase overall profitability. To explore this potential, this paper proposes an optimal investment portfolio model. First, we model and compare the active and reactive power characteristics of ELZs powered by TRs and IGBT-Rs. Second, we consider the investment of ELZs, rectifiers, and var resources and coordinate the operation of renewables, energy storage, var resources, and the on-off switching and load allocation of multiple ELZs. Subsequently, a two-stage stochastic programming (SP) model based on weighted information gap decision theory (W-IGDT) is developed to address the uncertainties of the renewable power and hydrogen price, and we apply the progressive hedging (PH) algorithm to accelerate its solution. Case studies demonstrate that optimal rectifier configurations increase revenue by at most 13.78% compared with configurations using only TRs or IGBT-Rs, existing project setups, or intuitive designs. Under the optimal portfolio, reactive power compensation investment is nearly eliminated, with a preferred TR-to-IGBT-R ratio of 3:1.
Nomenclature
@@
Active power of the ELZ/rectifier
The active loss of the TR/IGBT-R of the ELZ
@@ I-A Background and Motivation
Renewable power-to-hydrogen (ReP2H) facilitates wind and solar integration [1] and has significant potential for applications in energy storage, transportation [1], and the chemical industry [2]. With the continuous growth of renewable energy (RE) installations, the ReP2H industry has grown rapidly under policy incentives and market drivers [3, 4], leading to increased investment in electrolyzers (ELZs) in large-scale hydrogen plants to replace fossil-based hydrogen production.
Industrial hydrogen plants typically consist of multiple ELZs [5], each integrating a rectifier, an electrolytic stack, and the balance of plant (BoP) [6]. Since electrolytic stacks operate on DC power, rectifiers are required to convert AC to DC. Currently, two types of rectifiers are commercially available and widely used in engineering: thyristor rectifiers (TRs) and voltage-source insulated-gate bipolar transistor-based rectifiers (IGBT-Rs) [7, 8, 9]. They offer distinct tradeoffs, as summarized in Table I. Besides these, diode-based rectifiers (DRs), current-source rectifiers (CSRs) [10], and hybrid thyristor-IGBT rectifiers (HRs) [11] have also been studied. However, due to the limited controllability of DRs, the immaturity of the CSRs [10], and the engineering complexity of HRs, these rectifiers are not commercially available on large scales. Hence, only SCRs and IGBT-Rs are adopted in utility-scale ReP2H projects, and this work focuses on these two types.
IGBT-Rs use PWM modulation and voltage vector control, enabling independent active and reactive power control. Thus, IGBT-Rs can maintain a power factor (PF) of 1 [7] or provide reactive power support [12], or even enable low voltage ride-through [13]. Moreover, their fast response makes IGBT-Rs well suited for weak or off-grid systems [14]. However, their structural complexity leads to higher costs and power losses.
@@
In practice, some large-scale projects [8, 9] use mixed rectifier (MR) configurations to leverage the complementary characteristics of TRs and IGBT-Rs. The China Energy Engineering Songyuan Hydrogen Industry Park [8] in Songyuan, China, employs 32 TRs and 32 IGBT-Rs to power 64 alkaline ELZs, and the Wind-Solar Hydrogen-Ammonia Integration Project [9] in Da’an, China, uses 12 TRs and 24 IGBT-Rs for 36 ELZs. However, to the best of the author’s knowledge, such configurations have not been quantitatively analyzed. Overlooking rectifier performance in the planning stage may hinder coordination between the hydrogen plant and the electrical network, as the impact of reactive power loads and resources on ReP2H system economics becomes hard to quantify. Suboptimal configurations may lead to excessive or insufficient reactive power compensation and reduced hydrogen yield, increasing investment or operating costs. Illustrative examples are presented in Section II-D.
Specifically, the optimal investment of rectifiers in a ReP2H system is a comprehensive optimization problem rather than a simple pairing of ELZs with rectifiers. It needs multiple interrelated considerations:
- 1.
2.
Coupling between active and reactive power flows in the electrical network of the ReP2H system.
3.
Coupling between the hydrogen production process and the network power flow.
However, existing ReP2H planning research has focused mainly on the sizing of energy sources and ELZs while considering only active power and neglecting rectifier selection, the electrical network, and var resources. Coordinating investment planning and production scheduling with respect to both active and reactive power is vital for improving capital efficiency, which, unfortunately, has not been investigated.
@@ I-B Literature Review and Research Gap
ReP2H planning research typically addresses RE, battery energy storage (BES), ELZs, and hydrogen storage (HS). For example, Li et al. [19] proposed an investment model for ELZs and seasonal HS, considering power and hydrogen delivery, to address the spatiotemporal imbalance between renewable resources and hydrogen demand. Pan et al. [20] developed a robust capacity sizing model for wind turbines (WTs), photovoltaics (PVs), ELZs, and HSs in integrated energy systems. Zhu et al. [21] proposed an energy management framework and studied BES capacity sizing in off-grid ReP2H systems. Li et al. [22] explored the configuration of the number and size of ELZs and developed operation rules to enhance wind power integration. To assess the cost competitiveness of green hydrogen, Ibáñez-Rioja et al. [23] optimized planning and operation using 30 years of real-world wind and solar data. Similarly, Zheng et al. [24] analyzed the levelized cost of hydrogen (LCOH) for off-grid wind P2H systems and reported that optimal sizing reduces LCOH. However, the aforementioned studies [19, 20, 21, 22, 23, 24] overlooked rectifier configuration and reactive power balance in ReP2H planning. This omission could potentially cause ineffective var resource sizing from disregarded rectifier reactive characteristics and reduce hydrogen productivity due to unconsidered rectifier losses. Moreover, given that rectifiers account for 20–30% of ELZ investment [25], their optimal selection is crucial for economic feasibility.
The effectiveness of planning is closely associated with operational performance. Regarding P2H scheduling, Wang et al. [26] proposed a two-phase control strategy that considers the minimum power constraint of ELZs, determining the on-off switching and power to maximize RE utilization while balancing their lifespan. Liang et al. [27] designed an ELZ scheduling strategy using the Pelican Optimization Algorithm to improve P2H efficiency. Firdous et al. [28] developed a modular P2H management model in integrated power-to-ammonia systems to enhance operation flexibility. Khajeh et al. [29] proposed a wind-hydrogen system scheduling scheme to provide frequency support and congestion management for the transmission system, thereby enhancing its profitability. Han et al. [30] proposed a robust optimization framework for an integrated electricity-heat-hydrogen system that accounts for ELZ-network interactions. Even with their contributions, these studies did not account for rectifier reactive power or its coupling with the electrical network, making it difficult to quantify the impact of var resources on ReP2H system economics. Based on rectifier performance, the active-reactive power coordination between hydrogen production and the network is particularly important.
Current research on electrolysis rectifiers has focused mainly on performance analysis [15, 7] or short-term scheduling [6, 16]. For example, Ruuskanen et al. [15] investigated the power quality and reactive power characteristics of ELZs powered by TRs. Koponen et al. [7] compared the tradeoffs of the TR and IGBT-R in terms of performance and investment in industrial applications. In our prior work [16, 6], Li et al. [16] integrated rectifier power characteristics into ReP2H scheduling, revealing conflicts between productivity and the grid-side PF in grid-connected hydrogen plants, whereas Zeng et al. [6] addressed the tradeoffs between P2H energy conversion efficiency and network losses caused by the reactive power of a TR and proposed a coordinated power management approach for large hydrogen plants that integrate multiple ELZs.
@@ I-C Contributions of This Work
The active and reactive characteristics of TR- and IGBT-R-powered ELZs are modeled and compared, demonstrating their complementarity. Further analysis shows that an MR configuration enables more effective coordination between hydrogen production and network-side performance in ReP2H systems, as compared to strategies focused solely on hydrogen plant operation.
@@
An optimal investment portfolio model for MR and var resources in ReP2H systems is proposed, in which the operation stage adopts a coordinated active-reactive power scheduling strategy for hydrogen production and electrical network. The model handles uncertainties via weighted information gap decision theory (W-IGDT). A two-stage decomposition algorithm, progressive hedging (PH) [31], is applied to improve computational efficiency.
@@
Case studies on realistic systems show that optimizing MR configuration can increase revenue by at most 13.78% compared with that of uniform configurations or those in existing projects and intuitive designs. The results suggest that a 3:1 ratio of TRs to IGBT-Rs offers better benefits and is expected to provide a reference for industrial practices.
@@ II-A General Configuration of ReP2H Systems
Based on existing projects [8, 9], the general structure of a ReP2H system is illustrated in Fig. 1. The system integrates RE sources (wind and solar), transmission lines, a hydrogen plant, and sometimes a BES. The hydrogen plant often comprises multiple ELZs. Due to policy restrictions [32] and PF constraints [16], which aim to ensure local renewable utilization, facilitate green hydrogen certification, and avoid occupying the transmission capacity, both active and reactive power exchange with the external grid is limited. Therefore, such exchanges are not considered in this study.
@@ II-C Reactive Power Characteristics
where is the turn ratio of the rectifier transformer and where is the harmonic factor of the 24-pulse TR. The detailed derivation can be found in [6].
Since and are functions of and , substituting (3) and (5) into (18) yields the complete expression for , which is employed for MR configuration planning in Section III. As the OLTC can regulate the AC voltage at the hydrogen plant, thereby affecting the TR’s firing angle and its reactive power demand, coordinated control of the OLTC and TR can be used to optimize system-level operation. The detailed formulation is presented in Table III and discussed in Section III.
@@ II-C2 IGBT-R
II-D Performance Comparison of Different MR Configurations
This section compares different examples to show that intuitive rectifier configurations may fall short, while an MR setup offers better performance.
II-D1 A single rectifier
@@
II-D2 MR configuration
To illustrate the impact of different rectifier configurations, three examples are given. In the first example, a hydrogen plant with four 5 MW ELZs supplied with 12 MW of power for one hour. Using hydrogen production maximization as the objective, we calculate the hydrogen yield and reactive demand for each configuration. Fig. 5(a) presents the results along with the investment costs. TRs offer lower costs and slightly higher hydrogen yield. If the interaction between hydrogen production and the electrical network is omitted, the all-TR setup seems the optimal choice. However, as discussed in Section II-A, both voltage security and network losses must be carefully addressed. Thus, further analysis is carried out based on a simple illustrative system, as shown in Fig. 6.
In the second example, the power source remains at 12 MW. However, due to network losses, the actual power delivered to the hydrogen plant is reduced. The optimized performance under different configurations is presented in Fig. 5(b). In contrast to the first case, the all-TR setup results in the lowest hydrogen production, primarily due to significant network losses induced by high reactive power demand. Since IGBT-Rs are capable of providing reactive power support, incorporating them enhances overall system performance. When considering both investment cost and hydrogen yield, the 4-TR setup is no longer optimal. An MR configuration improves hydrogen yield with a modest cost increase, and the 3 TR + 1 IGBT-R configuration achieves the best performance.
In the third example, voltage security and network losses are explicitly considered, making var compensation essential. To isolate its impact, we assume full compensation for the hydrogen plant’s reactive load. The corresponding results are shown in Fig. 5(c). Although the 4-TR configuration delivers the highest productivity, the additional cost of the var compensator increases the total investment. In comparison, using one IGBT-R achieves similarly good productivity with lower total cost. Together with the second example, this confirms the necessity of coordinating hydrogen production with the electrical network, offering insights that challenge intuitive preferences.
It should be noted, however, that these examples represent only a single-time snapshot. Whether an MR configuration can optimally balance these factors over a longer continuous operation period requires further investigation.
@@ III Optimal Investment Portfolio Planning of Mixed Rectifier Configurations
III-A Planning and Operational Framework
The optimal investment portfolio of MRs is a mixed planning and operation problem, consisting of two stages and their interaction:
@@ III-A Planning and Operational Framework
In the framework summarized in Fig. 7, the investment decisions are passed to the operation stage once determined. The operation stage then performs simulations with the objective of maximizing revenue and feeds the result back to the investment stage, similar to the procedure adopted in [36].
@@ III-B Objective Function
Here, and are the discount rate and equipment lifespan, respectively; represents the probability of scenario ; and is the annual O&M cost factor. In (25), maximizing hydrogen yield simultaneously achieves maximum RE utilization and reactive power optimization, while also accounting for SVG planning and network losses.
@@ III-C2 Operational constraints
We consider the coordination among active and reactive resources such as the electrical network, ELZs, WTs, PV plants, BES, var compensation, and OLTC. Their models are presented in Table III, where we use the DistFlow model [41] to describe the network power flow, which is radial in common ReP2H systems.
In operation, active power balance is achieved by on-off switching and load allocation among multiple ELZs in the hydrogen plant, assisted by the BES. The reactive power flow is optimized in coordination with the active power to support the voltage and reduce network losses. For details of the operation model, readers are referred to [6].
@@ III-D2 W-IGDT-based portfolio model
Although the investment decisions deal with a long-term time frame considering the lifespan of equipment, the optimization problem incorporates a short-term scheduling program, in which the operator needs to handle the intraday uncertainty of renewable power. To address the short-term uncertainty, a risk-averse model based on robust IGDT [39] is developed. Based on the revenue of (46), with a preset deviation factor , the robust IGDT model maximizes the uncertainty boundary under the worst-case renewable generation, as follows:
@@ III-E Two-stage Decomposition-Based Solution Approach—PH
The overall solution procedure of the proposed method and its application to MR configuration and var resource planning in the ReP2H system are summarized in Fig. 8.
@@ IV Case Studies
Case studies are performed based on two real-world systems in Inner Mongolia, China, as shown in Fig. 9. The test systems are adapted from our prior work [6]. The small-scale system validates the proposed portfolio model and explores the potential of MR configurations, whereas the large-scale system verifies the scalability of the model and provides insights for practice. Simulations are performed via Wolfram Mathematica 14.0, with optimizations solved via Gurobi 11.0.0.
@@ IV-A2 Large-scale system
| ELZ w/o rectifier | 5 MW | 7,000,000 CNY [44] | 8% | 20 years |
| TR | 6 MVA | 1,000,000 CNY [17] | ||
| IGBT-R | 6 MVA | 2,200,000 CNY [17] | ||
| SVG | 350 CNY/kVar [45] |
| Parameter | Value | Parameter | Value |
| 7.8J/℃ | 1.06 ℃/kW | ||
| 17 kW/℃ | / | 5 ℃ / 4.0 | |
| / | 12 / 2.4 kA | / | 80 / 25 ℃ |
| 0.5 MW | / | 1000 / 0 CNY | |
| / | 0.9 / 0.1 p.u. | 0.5 / 0.5 p.u. | |
| 0.95 / 0.95 | / | 1.05 / 0.95 p.u. | |
| (35kV) | 0.37 p.u. | / (35kV) | 0.0096 / 0.0262 p.u./km |
| (220kV) | 1.8 p.u. | / (220kV) | 0.00025 / 0.000832 p.u./km |
The key investment parameters and operation-related parameters are listed in Tables IV and V, respectively. More details on the topology and operational details of the ReP2H system and ELZs are available in [6]. Renewable generation for typical days, derived from on-site historical data via K-means clustering, is shown in Figs. 10(a) and 10(b). The hydrogen price scenario is shown in Fig. 10(c), with fluctuations modeled based on local agricultural ammonia demand [46].